OpenStudy (anonymous):

f(x)=sqrt x+2, how do you use differential for this, dy/dx?

OpenStudy (bradely):

the function is (sqrt x)+2 is it correct

OpenStudy (bradely):

or sqrt(x+2)

OpenStudy (anonymous):

its sqrt(x=2)

OpenStudy (bradely):

step by step solution http://www.mathskey.com/question2answer/5551/how-to-do-this?show=5552#a5552

OpenStudy (anonymous):

actually im not sure how did u do this starting from this part =1/(2sqrt(x+1))*1 =1/(2sqrt(x+1))

OpenStudy (bradely):

Using formula d/dx(x^n)=x^(n-1) d/dx(sqrt(x+1))=(1/2)(x+1)^(1/2-1)*d/dx(x+1) =(1/2)(x+1)^(-1/2)*1 =1/(2sqrt (x+1))

OpenStudy (anonymous):

alright Bradely.. Thank u so much. u have been very very helpful..Now i got it! haha :D